OpenStudy (anonymous):

Write the expression in terms of sinx and cosx only: cos^2(2x)-sin(2x) Please help! This is on the review in our PreCalc textbook (high school junior, please make it so that I can understand, thanks) but there isn't an example in the book that I can understand! Thanks for the help :) By the way, it says the answer should be 1-4sin^2xcos^2x-2sinxcosx.

OpenStudy (anonymous):

I don't even know anymore. I don't know what the book wants me to do. I've tried doing it, but I can't get the answer it wants me to... :/ Thanks anyway.

OpenStudy (superdavesuper):

u know sin(2x) = 2 sin(x)cos(x) right?

OpenStudy (beccaboo333):

Sorry I do not know

OpenStudy (superdavesuper):

use the following equations to solve this prob: sin(2x) = 2sin(x)cos(x) , cos(2x) = cos^2(x) - sin^2(x) and cos^2(2x) = cos(2x)*cos(2x) it will take a few multiplication n simplification but u will get the right ans. good luck!

OpenStudy (anonymous):

y = cos^2 (2x) - sin (2x) = (cos^2 x - sin^2 x)^2 - 2sin x*cos x = = cos^4 x + sin^4 x - 2(cos^2 x)*(sin^2 x) - 2sin x*cos x. (1). Find the value of the sum (sin^4 x + cos^4 x) (sin^4x + cos^4 x) = (sin^2 x + cos^2 x)^2 = 1 = = sin^4 x + cos^4 x + 2sin^2 x*cos^2 x. sin^4 x + cos^4 x = 1 - 2sin^2 x*cos^2 x (2) Substitute this value (2) into equation (1), we get: y = 1 - 4sin^2 x*cos^2* x - 2sin x*cos x